OpenStudy (anonymous):

Andrew Goes For A Walk Every Night After Work. A Mop Of The Route Has A Scale Of 1:30,000 In CM If Andrew Walks 1.8 Kilometers On Monday Night, How Many CM Is The Route On The Map ?

OpenStudy (anonymous):

Your conversion rate is 1:30,000. So for every 1Km, there is (1/30,000)Cm. If there's 1.8Km, how would you find the respective scale?

hero (hero):

For it to make sense, the scale should have said 1Km = 30,000 cm

hero (hero):

Then the next step would be to set up proportion: \[\frac{1 \space \text{km}}{1.8 \space \text{km}} = \frac{30000 \space \text{cm}}{x \space \text{cm}}\] Then cancel the km and cm to get: \[\frac{1}{1.8} = \frac{30000}{x}\] Then cross multiply to get x = 30000(1.8) x = ?

OpenStudy (anonymous):

54000

hero (hero):

Yup